# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v2

@@ C.2 Proof of Lemma 8
$\left|\Braket{\phi}{f(1)-\left(f(0)+f^{\prime}(0)+\frac{1}{2}f^{\prime\prime}(0)\right)}{\phi}\right|$ | $\leq\left\|f(1)-\left(f(0)+f^{\prime}(0)+\frac{1}{2}f^{\prime\prime}(0)\right)\right\|_{1}$ (218)
@@
$\braket{\phi}{O^{\prime n}O^{m}}{\phi}=\begin{cases}1&\quad(n,m)=(0,0)\\ \braket{\phi}{O}{\phi}&\quad(n,m)=(0,1)\\ 0&\quad(n,m)=(1,0)\end{cases},$ (220)
@@
$\braket{\phi}{f(0)}{\phi}$ $=1,$ (221)
[…]
@@
$\braket{\phi}{O^{\prime n}O^{m}}{\phi}=\begin{cases}\braket{\phi}{O^{2}}{\phi}&\quad(n,m)=(0,2)\\ \braket{\phi}{O(I-\phi)O}{\phi}&\quad(n,m)=(1,1),\,(2,0)\end{cases}$ (223)
@@
[…]
@@
$\braket{\phi}{f^{\prime\prime}(0)}{\phi}$ $=\sum_{\begin{subarray}{c}n,m,n^{\prime},m^{\prime}=0,1,2\\ n+m+n^{\prime}+m^{\prime}=2\end{subarray}}\frac{(-\mathrm{i})^{n}}{n!}\frac{\mathrm{i}^{m}}{m!}\frac{(-\mathrm{i})^{m^{\prime}}}{m^{\prime}!}\frac{\mathrm{i}^{n^{\prime}}}{n^{\prime}!}\braket{\phi}{O^{\prime n}O^{m}}{\phi}\braket{\phi}{O^{m^{\prime}}O^{\prime n^{\prime}}}{\phi}$ (225)
[…]
Therefore, by using $\braket{\phi}{f(1)}{\phi}=\mathrm{Fid}\left(\phi_{\bm{u}/\sqrt{N}},\tilde{\phi}_{\bm{u}/\sqrt{N}}\right)$, Eq. (219) implies
@@
Therefore, as long as the asymptotic properties of an i.i.d. copies of $\phi_{\bm{u}/\sqrt{N}}=e^{\mathrm{i}\bm{u}\cdot\bm{X}/\sqrt{N}}\phi e^{-\mathrm{i}\bm{u}\cdot\bm{X}/\sqrt{N}}$ is concerned, it suffice to analyze $\tilde{\phi}_{\bm{u}/\sqrt{N}}\coloneqq e^{\mathrm{i}\bm{u}\cdot\tilde{\bm{X}}/\sqrt{N}}\phi e^{\mathrm{i}\bm{u}\cdot\tilde{\bm{X}}/\sqrt{N}}$.
@@
Let |$\ket{\phi}$ be an arbitrary state in a $d$-dimensional Hilbert space. For any fixed set of Hermitian operators $\{X_{i}\}_{i=1}^{m}$, we define a pure-state qudit model $\tilde{\phi}_{\bm{\theta}}$ by
|$\ket{\tilde{\phi}_{\bm{\theta}}}$ $\coloneqq\exp\left(-\mathrm{i}\sum_{i=1}^{m}\theta^{i}\tilde{X}_{i}\right)\ket{\phi},\quad\tilde{X}_{i}\coloneqq(I-\phi)X_{i}\phi+\phi X_{i}(I-\phi).$ (231)
@@
$\mathrm{i}C_{ki}\coloneqq\braket{k}{\tilde{X}_{i}}{0},$ (232)
where {|$\{\ket{k}\}_{k=1}^{d-1}$ denotes an orthonormal basis of the subspace orthogonal to |$\ket{\phi}$.
Let $\mathcal{H}_{d-1}$ denote the $(d-1)$-mode Fock space. For $\bm{z}=(z_{1},\cdots,z_{d-1})^{\top}\in\mathbb{C}^{d-1}$, we denote the coherent state by
|$\ket{\bm{z}}\coloneqq\left(\bigotimes_{i=1}^{d-1}e^{z_{i}a_{i}^{\dagger}-z_{i}^{*}a_{i}}\ket{0}_{i}\right)\in\mathcal{H}_{d-1},$ (233)
where $a_{i}^{\dagger},a_{i}$ and |$\ket{0}_{i}$ denote the creation and annihilation operator and the vacuum state for the $i$th mode.
@@
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{T}_{N}(\tilde{\phi}^{\otimes N}_{\bm{u}/\sqrt{N}}),\ket{C\bm{u}}\bra{C\bm{u}}\right)$ $=0,$ (234)
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\tilde{\phi}^{\otimes N}_{\bm{u}/\sqrt{N}},\mathcal{S}_{N}(\ket{C\bm{u}}\bra{C\bm{u}})\right)$ $=0.$ (235)
@@
Let |$\ket{\phi}$ be an arbitrary pure state in a finite-dimensional Hilbert space $\mathcal{H}$. Fix any sets of Hermitian operators $\bm{X}=\{X_{i}\}_{i=1}^{m}$ and introduce a pure-state statistical model $\phi_{\bm{\theta}}\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}}\phi e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}}$ for $\bm{\theta}\in\mathbb{R}^{m}$. For any $r>0$, there exist sequences of quantum channels $\{\mathcal{E}_{N}\}_{N}$ and $\{\mathcal{E}^{\prime}_{N}\}_{N}$ such that
@@
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<\lfloor rN\rfloor^{\epsilon}}T\left(\mathcal{T}_{\lfloor rN\rfloor}\left(\tilde{\phi}^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{\lfloor rN\rfloor}}\right),\ket{C\bm{u}}\bra{C\bm{u}}\right)$ $=0,$ (238)
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<{\lfloor rN\rfloor}^{\epsilon}}T\left(\tilde{\phi}^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{\lfloor rN\rfloor}},\mathcal{S}_{\lfloor rN\rfloor}(\ket{C\bm{u}}\bra{C\bm{u}})\right)$ $=0.$ (239)
@@
$\lim_{N\to\infty}\sup_{\|\bm{u}^{\prime}\|<\lfloor rN\rfloor^{\epsilon}\sqrt{\frac{\lfloor rN\rfloor}{N}}}T\left(\mathcal{T}_{\lfloor rN\rfloor}\left(\tilde{\phi}^{\otimes\lfloor rN\rfloor}_{\bm{u}^{\prime}/\sqrt{N}}\right),\Ket{\sqrt{\frac{\lfloor rN\rfloor}{N}}C\bm{u}^{\prime}}\Bra{\sqrt{\frac{\lfloor rN\rfloor}{N}}C\bm{u}^{\prime}}\right)$ $=0,$ (240)
$\lim_{N\to\infty}\sup_{\|\bm{u}^{\prime}\|<{\lfloor rN\rfloor}^{\epsilon}\sqrt{\frac{\lfloor rN\rfloor}{N}}}T\left(\tilde{\phi}^{\otimes\lfloor rN\rfloor}_{\bm{u}^{\prime}/\sqrt{N}},\mathcal{S}_{\lfloor rN\rfloor}\left(\Ket{\sqrt{\frac{\lfloor rN\rfloor}{N}}C\bm{u}^{\prime}}\Bra{\sqrt{\frac{\lfloor rN\rfloor}{N}}C\bm{u}^{\prime}}\right)\right)$ $=0.$ (241)
@@
$\lim_{N\to\infty}\left|\Braket{\sqrt{\frac{\lfloor rN\rfloor}{N}}C\bm{u}^{\prime}}{\sqrt{r}C\bm{u}^{\prime}}\right|=\lim_{N\to\infty}e^{-\|C\bm{u}^{\prime}\|^{2}\left(\sqrt{r}-\sqrt{\frac{\lfloor rN\rfloor}{N}}\right)^{2}}=1$ (243)
@@
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{T}_{\lfloor rN\rfloor}\left(\tilde{\phi}^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{N}}\right),\ket{\sqrt{r}C\bm{u}}\bra{\sqrt{r}C\bm{u}}\right)$ $=0,$ (244)
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon^{\prime}}}T\left(\tilde{\phi}^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{N}},\mathcal{S}_{\lfloor rN\rfloor}\left(\ket{\sqrt{r}C\bm{u}}\bra{\sqrt{r}C\bm{u}}\right)\right)$ $=0.$ (245)
@@
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{T}_{N}\left(\tilde{\phi}_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N}\right),\ket{\sqrt{r}C\bm{u}}\bra{\sqrt{r}C\bm{u}}\right)$ $=0,$ (246)
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\tilde{\phi}_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N},\mathcal{S}_{N}(\ket{\sqrt{r}C\bm{u}}\bra{\sqrt{r}C\bm{u}})\right)$ $=0.$ (247)
@@
Importantly, for both $\phi_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N}$ and $\phi^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{N}}$, the QGT with respect to the parameters $\bm{u}$ is equal to $r\,\mathrm{Cov}(\phi,\bm{X})$ in the limit of $N\to\infty$, where the elements of covariance matrix are defined by $\left(\mathrm{Cov}(\phi,\bm{X})\right)_{ij}\coloneqq\braket{\phi}{X_{i}(I-\phi)X_{j}}{\phi}$. In other words, the QGTs are preserved under reversible asymptotic conversion between $\{\phi_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N}\}_{N}$ and $\{\phi^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{N}}\}_{N}$.
From the above lemma states, the asymptotic rate can be effectively set to $r=1$ by adjusting the scale of the parameters. The following lemma claims that the asymptotic conversion with rate $r=1$ is possible only if the covariance matrix of the original state is larger than that of the target state in the sense of matrix inequality:
@@
Let |$\ket{\psi}\in\mathcal{H}$ and |ϕketitalic-$\ket{\phi}\in\mathcal{H}^{\prime}$ be arbitrary pure states in finite-dimensional Hilbert spaces $\mathcal{H},\mathcal{H}^{\prime}$. Fix any sets of Hermitian operators $\bm{X}=\{X_{i}\}_{i=1}^{m}$ and $\bm{X}^{\prime}=\{X_{i}^{\prime}\}_{i=1}^{m}$. If $\mathrm{Cov}(\psi,\bm{X})\geq\mathrm{Cov}(\phi,\bm{X}^{\prime})$ holds, then there exists a quantum channel $\mathcal{E}$ such that
@@
Let us fix orthonormal bases {|$\{\ket{k}\}_{k=1}^{d-1}$ and {|$\{\ket{k^{\prime}}\}_{k^{\prime}=1}^{d^{\prime}-1}$ for the orthogonal complements to |$\ket{\psi}$ and |$\ket{\phi}$, respectively, where $d\coloneqq\dim\mathcal{H}$ and $d^{\prime}\coloneqq\dim\mathcal{H}^{\prime}$. For matrices $C:\mathbb{C}^{m}\to\mathbb{C}^{d-1}$, $C^{\prime}:\mathbb{C}^{m}\to\mathbb{C}^{d^{\prime}-1}$ whose elements are given by
@@
$K_{k}$ $\coloneqq\sum_{l=1}^{d-1}\left(\sqrt{\Gamma}\right)_{kl}\ket{\phi}\bra{l}\qquad(k=1,2,\cdots,d-1),$ (260)
@@
$\sum_{k=0}^{d-1}K_{k}^{\dagger}K_{k}$ =|$=\ket{\psi}\bra{\psi}+\sum_{k,l=1}^{d-1}\sum_{k^{\prime},l^{\prime}=1}^{d^{\prime}-1}z_{k^{\prime}k}^{*}z_{l^{\prime}l}\delta_{k^{\prime}l^{\prime}}\ket{k}\bra{l}+\sum_{k,l=1}^{d-1}\sum_{m=1}^{d-1}\left(\sqrt{\Gamma}\right)_{mk}^{*}\left(\sqrt{\Gamma}\right)_{ml}\ket{k}\bra{l}$ | (261)
@@
$\left|\Braket{\phi}{f(1)-\left(f(0)+f^{\prime}(0)+\frac{1}{2}f^{\prime\prime}(0)\right)}{\phi}\right|$ | $\leq\left\|f(1)-\left(f(0)+f^{\prime}(0)+\frac{1}{2}f^{\prime\prime}(0)\right)\right\|_{1}$ (264)
@@
Let us evaluate each term of $\braket{\phi}{\left(f(0)+f^{\prime}(0)+\frac{1}{2}f^{\prime\prime}(0)\right)}{\phi}$. From the definition of the channel $\mathcal{E}$, we get $\mathcal{E}(\psi)=\phi$, implying that
$\braket{\phi}{f(0)}{\phi}=1.$ (266)
Since $\braket{\phi}{O^{\prime}}{\phi}=0$, we have $\braket{\phi}{O^{\prime}\mathcal{E}(\psi)}{\phi}=0$ and $\braket{\phi}{\mathcal{E}(\psi)O^{\prime}}{\phi}=0$. In addition, since
$\mathcal{E}(O\psi)$ $=\frac{-\mathrm{i}}{\sqrt{N}}\sum_{k^{\prime}=1}^{d^{\prime}-1}\left(ZC\bm{u}\right)_{k^{\prime}}\ket{k^{\prime}}\bra{\phi},$ (267)
we also get $\braket{\phi}{\mathcal{E}(O\psi)}{\phi}=0$ and $\braket{\phi}{\mathcal{E}(\psi O)}{\phi}=0$. Therefore,
$\braket{\phi}{f^{\prime}(0)}{\phi}=0.$ (268)
From $\mathcal{E}(\psi)=\phi$ and $\braket{\phi}{O^{\prime}}{\phi}=0$, we get
$\frac{1}{2}\braket{\phi}{(O^{\prime 2}\mathcal{E}(\psi)+\mathcal{E}(\psi)O^{\prime 2})}{\phi}$ $=\braket{\phi}{O^{\prime 2}}{\phi}=\frac{1}{N}\sum_{i,j=1}^{m}u^{i}u^{j}\braket{\phi}{\tilde{X}^{\prime}_{i}\tilde{X}^{\prime}_{j}}{\phi}=\frac{1}{N}(C^{\prime{\dagger}}\bm{u})^{\dagger}C^{\prime}\bm{u},$ (269)
$\braket{\phi}{O^{\prime}\mathcal{E}(\psi)O^{\prime}}{\phi}$ $=0.$ (270)
From Eq. (267), we find $\bra{\phi}\mathcal{E}(O\psi)=0$ and $\mathcal{E}(\psi O)\ket{\phi}=0$. Therefore,
$\braket{\phi}{\mathcal{E}(O\psi)O^{\prime}}{\phi}=0,\quad\braket{\phi}{O^{\prime}\mathcal{E}(\psi O)}{\phi}=0.$ (271)
@@
$O^{\prime}\mathcal{E}(O\psi)=\frac{1}{N}(C^{\prime}\bm{u})^{\dagger}ZC\bm{u}\ket{\phi}\bra{\phi},\quad\mathcal{E}(\psi O)O^{\prime}=\frac{1}{N}(ZC\bm{u})^{\dagger}C^{\prime}\bm{u}\ket{\phi}\bra{\phi},$ (272)
@@
$\braket{\phi}{O^{\prime}\mathcal{E}(O\psi)}{\phi}$ $=\frac{1}{N}(C^{\prime}\bm{u})^{\dagger}ZC\bm{u},\quad\braket{\phi}{\mathcal{E}(\psi O)O^{\prime}}{\phi}=\frac{1}{N}(ZC\bm{u})^{\dagger}C^{\prime}\bm{u}.$ (273)
@@
$\mathcal{E}(O^{2}\psi)$ $=\frac{1}{N}\mathcal{E}((C\bm{u})^{\dagger}(C\bm{u})\ket{\psi}\bra{\psi})=\frac{1}{N}(C\bm{u})^{\dagger}(C\bm{u})\ket{\phi}\bra{\phi},$ (274)
$\mathcal{E}(O\psi O)$ $=\frac{1}{N}\sum_{k,l=1}^{d-1}(C\bm{u})_{k}(C\bm{u})_{l}^{*}\mathcal{E}\left(\ket{k}\bra{l}\right)=\frac{1}{N}\left(\sum_{k^{\prime},l^{\prime}=1}^{d^{\prime}-1}(ZC\bm{u})_{k^{\prime}}(ZC\bm{u})_{l^{\prime}}^{*}\ket{k^{\prime}}\bra{l^{\prime}}+(C\bm{u})^{\dagger}\Gamma C\bm{u}\ket{\phi}\bra{\phi}\right),$ (275)
@@
$\frac{1}{2}\braket{\phi}{\mathcal{E}(O^{2}\psi+\psi O^{2})}{\phi}$ $=\frac{1}{N}(C\bm{u})^{\dagger}C\bm{u},\quad\braket{\phi}{\mathcal{E}(O\psi O)}{\phi}=\frac{1}{N}(C\bm{u})^{\dagger}\Gamma(C\bm{u}).$ (276)
@@
$\Braket{\phi}{\frac{1}{2}f^{\prime}(0)}{\phi}$ $=-\frac{1}{2}\braket{\phi}{(O^{\prime 2}\mathcal{E}(\psi)-2O^{\prime}\mathcal{E}(\psi)O^{\prime}+\mathcal{E}(\psi)O^{\prime 2})}{\phi}+\braket{\phi}{(O^{\prime}\mathcal{E}(O\psi-\psi O)-\mathcal{E}(O\psi-\psi O)O^{\prime})}{\phi}$
$\quad-\frac{1}{2}\braket{\phi}{\mathcal{E}(O^{2}\psi-2O\psi O+\psi O^{2})}{\phi}$ (277)
@@
$\Braket{\phi}{\frac{1}{2}f^{\prime}(0)}{\phi}=0.$ (281)
Since $\mathrm{Fid}\left(\mathcal{E}(\tilde{\psi}_{\bm{u}/\sqrt{N}}),\tilde{\phi}_{\bm{u}/\sqrt{N}}\right)=\braket{\phi}{f(1)}{\phi}$, from Eqs. (265), (266), (268), and (281), we have proven

### v3

@@ C.2 Proof of Lemma 8
$\left|\Braket{\phi|f(1)-\left(f(0)+f^{\prime}(0)+\frac{1}{2}f^{\prime\prime}(0)\right)|\phi}\right|$ $\leq\left\|f(1)-\left(f(0)+f^{\prime}(0)+\frac{1}{2}f^{\prime\prime}(0)\right)\right\|_{1}$ (262)
@@
$\braket{\phi|O^{\prime n}O^{m}|\phi}=\begin{cases}1&\quad(n,m)=(0,0)\\ \braket{\phi|O|\phi}&\quad(n,m)=(0,1)\\ 0&\quad(n,m)=(1,0)\end{cases},$ (264)
@@
$\braket{\phi|f(0)|\phi}$ $=1,$ (265)
$\braket{\phi|f^{\prime}(0)|\phi}$ $=\mathrm{i}\braket{\phi|O|\phi}-\mathrm{i}\braket{\phi|O|\phi}=0.$ (266)
@@
$\braket{\phi|O^{\prime n}O^{m}|\phi}=\begin{cases}\braket{\phi|O^{2}|\phi}&\quad(n,m)=(0,2)\\ \braket{\phi|O(I-\phi)O|\phi}&\quad(n,m)=(1,1),\,(2,0)\end{cases}$ (267)
@@
$\sum_{\begin{subarray}{c}n,m=0,1,2\\ n+m=2\end{subarray}}\frac{(-\mathrm{i})^{n}}{n!}\frac{\mathrm{i}^{m}}{m!}\braket{\phi|O^{\prime n}O^{m}|\phi}=-\frac{1}{2}\braket{\phi|O^{2}|\phi}+\frac{1}{2}\braket{\phi|O(I-\phi)O|\phi}=-\frac{1}{2}\braket{\phi|O|\phi}^{2}$ (268)
@@
$\braket{\phi|f^{\prime\prime}(0)|\phi}$ $=\sum_{\begin{subarray}{c}n,m,n^{\prime},m^{\prime}=0,1,2\\ n+m+n^{\prime}+m^{\prime}=2\end{subarray}}\frac{(-\mathrm{i})^{n}}{n!}\frac{\mathrm{i}^{m}}{m!}\frac{(-\mathrm{i})^{m^{\prime}}}{m^{\prime}!}\frac{\mathrm{i}^{n^{\prime}}}{n^{\prime}!}\braket{\phi|O^{\prime n}O^{m}|\phi}\braket{\phi|O^{m^{\prime}}O^{\prime n^{\prime}}|\phi}$ (269)
$=-\frac{1}{2}\braket{\phi|O|\phi}^{2}+\braket{\phi|O|\phi}^{2}-\frac{1}{2}\braket{\phi|O|\phi}^{2}=0.$ (270)
Therefore, by using $\braket{\phi|f(1)|\phi}=\mathrm{Fid}\left(\phi_{\bm{u}/\sqrt{N}},\tilde{\phi}_{\bm{u}/\sqrt{N}}\right)$, Eq. (263) implies
@@
Therefore, as long as the asymptotic properties of an i.i.d. copies of $\phi_{\bm{u}/\sqrt{N}}=e^{\mathrm{i}\bm{u}\cdot\bm{X}/\sqrt{N}}\phi e^{-\mathrm{i}\bm{u}\cdot\bm{X}/\sqrt{N}}$ are concerned, it suffice to analyze $\tilde{\phi}_{\bm{u}/\sqrt{N}}\coloneqq e^{\mathrm{i}\bm{u}\cdot\tilde{\bm{X}}/\sqrt{N}}\phi e^{\mathrm{i}\bm{u}\cdot\tilde{\bm{X}}/\sqrt{N}}$.
@@
[…]
@@
$\mathrm{i}C_{ki}\coloneqq\braket{k|\tilde{X}_{i}|0},$ (276)
[…]
Let $\mathcal{H}_{d-1}$ denote the (d1)(d-1)-mode Fock space. For $\bm{z}=(z_{1},\cdots,z_{d-1})^{\top}\in\mathbb{C}^{d-1}$, we denote the coherent state by
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\lim_{N\to\infty}\left|\Braket{\sqrt{\frac{\lfloor rN\rfloor}{N}}C\bm{u}^{\prime}|\sqrt{r}C\bm{u}^{\prime}}\right|=\lim_{N\to\infty}e^{-\|C\bm{u}^{\prime}\|^{2}\left(\sqrt{r}-\sqrt{\frac{\lfloor rN\rfloor}{N}}\right)^{2}}=1$ (287)
@@
[…]
@@
[…]
@@
Importantly, for both $\phi_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N}$ and $\phi^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{N}}$, the QGT with respect to the parameters $\bm{u}$ is equal to $r\,\mathrm{Cov}(\phi,\bm{X})$ in the limit of $N\to\infty$, where the elements of covariance matrix are defined by $\left(\mathrm{Cov}(\phi,\bm{X})\right)_{ij}\coloneqq\braket{\phi|X_{i}(I-\phi)X_{j}|\phi}$. In other words, the QGTs are preserved under reversible asymptotic conversion between $\{\phi_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N}\}_{N}$ and $\{\phi^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{N}}\}_{N}$.
Using the above lemma, the asymptotic rate can be effectively set to $r=1$ by adjusting the scale of the parameters. The following lemma claims that the asymptotic conversion with rate $r=1$ is possible only if the covariance matrix of the original state is larger than that of the target state in the sense of matrix inequality:
@@
[…]
@@
$(C)_{ki}$ $\coloneqq-\mathrm{i}\braket{k|X_{i}|\psi},\quad(C^{\prime})_{l^{\prime}i}\coloneqq-\mathrm{i}\braket{l^{\prime}|X_{i}^{\prime}|\phi},$ (298)
@@
[…]
@@
[…]
@@
$\left|\Braket{\phi|f(1)-\left(f(0)+f^{\prime}(0)+\frac{1}{2}f^{\prime\prime}(0)\right)|\phi}\right|$ $\leq\left\|f(1)-\left(f(0)+f^{\prime}(0)+\frac{1}{2}f^{\prime\prime}(0)\right)\right\|_{1}$ (308)
@@
Let us evaluate each term of $\braket{\phi|\left(f(0)+f^{\prime}(0)+\frac{1}{2}f^{\prime\prime}(0)\right)|\phi}$. From the definition of the channel $\mathcal{E}$, we get $\mathcal{E}(\psi)=\phi$, implying that
$\braket{\phi|f(0)|\phi}=1.$ (310)
Since $\braket{\phi|O^{\prime}|\phi}=0$, we have $\braket{\phi|O^{\prime}\mathcal{E}(\psi)|\phi}=0$ and $\braket{\phi|\mathcal{E}(\psi)O^{\prime}|\phi}=0$. In addition, since
$\mathcal{E}(O\psi)$ $|,\displaystyle=\frac{-\mathrm{i}}{\sqrt{N}}\sum_{k^{\prime}=1}^{d^{\prime}-1}\left(ZC\bm{u}\right)_{k^{\prime}}\ket{k^{\prime}}\bra{\phi}$, (311)
we also get $\braket{\phi|\mathcal{E}(O\psi)|\phi}=0$ and $\braket{\phi|\mathcal{E}(\psi O)|\phi}=0$. Therefore,
$\braket{\phi|f^{\prime}(0)|\phi}=0.$ (312)
From $\mathcal{E}(\psi)=\phi$ and $\braket{\phi|O^{\prime}|\phi}=0$, we get
$\frac{1}{2}\braket{\phi|(O^{\prime 2}\mathcal{E}(\psi)+\mathcal{E}(\psi)O^{\prime 2})|\phi}$ $=\braket{\phi|O^{\prime 2}|\phi}=\frac{1}{N}\sum_{i,j=1}^{m}u^{i}u^{j}\braket{\phi|\tilde{X}^{\prime}_{i}\tilde{X}^{\prime}_{j}|\phi}=\frac{1}{N}(C^{\prime{\dagger}}\bm{u})^{\dagger}C^{\prime}\bm{u},$ (313)
$\braket{\phi|O^{\prime}\mathcal{E}(\psi)O^{\prime}|\phi}$ $=0.$ (314)
[…]
$\braket{\phi|\mathcal{E}(O\psi)O^{\prime}|\phi}=0,\quad\braket{\phi|O^{\prime}\mathcal{E}(\psi O)|\phi}=0.$ (315)
@@
[…]
@@
$\braket{\phi|O^{\prime}\mathcal{E}(O\psi)|\phi}$ $=\frac{1}{N}(C^{\prime}\bm{u})^{\dagger}ZC\bm{u},\quad\braket{\phi|\mathcal{E}(\psi O)O^{\prime}|\phi}=\frac{1}{N}(ZC\bm{u})^{\dagger}C^{\prime}\bm{u}.$ (317)
@@
[…]
@@
$\frac{1}{2}\braket{\phi|\mathcal{E}(O^{2}\psi+\psi O^{2})|\phi}$ $=\frac{1}{N}(C\bm{u})^{\dagger}C\bm{u},\quad\braket{\phi|\mathcal{E}(O\psi O)|\phi}=\frac{1}{N}(C\bm{u})^{\dagger}\Gamma(C\bm{u}).$ (320)
@@
$\Braket{\phi|\frac{1}{2}f^{\prime\prime}(0)|\phi}$ $=-\frac{1}{2}\braket{\phi|(O^{\prime 2}\mathcal{E}(\psi)-2O^{\prime}\mathcal{E}(\psi)O^{\prime}+\mathcal{E}(\psi)O^{\prime 2})|\phi}+\braket{\phi|(O^{\prime}\mathcal{E}(O\psi-\psi O)-\mathcal{E}(O\psi-\psi O)O^{\prime})|\phi}$
$\quad-\frac{1}{2}\braket{\phi|\mathcal{E}(O^{2}\psi-2O\psi O+\psi O^{2})|\phi}$ (321)
@@
$\Braket{\phi|\frac{1}{2}f^{\prime\prime}(0)|\phi}=0.$ (325)
Since $\mathrm{Fid}\left(\mathcal{E}(\tilde{\psi}_{\bm{u}/\sqrt{N}}),\tilde{\phi}_{\bm{u}/\sqrt{N}}\right)=\braket{\phi|f(1)|\phi}$, from Eqs. (309), (310), (312), and (325), we have proven